OpenStudy (anonymous):

How to simplify the expression: cos(4x)cos(3x) - sin(4x)cos(3x) Please guide me through, not the answer:)

OpenStudy (anonymous):

To simplify this expression, Am I looking for cos(theta +beta)?

OpenStudy (experimentx):

yes ... cos(4x+3x)

OpenStudy (anonymous):

ok, Is the answer to this a number? or is it something like the previous question?

OpenStudy (phi):

cos(4x+3x)= cos(4x)cos(3x)- sin(4x)sin(3x) It would be nice. Are we sure of the question?

OpenStudy (anonymous):

Is the answer 0.6087?

OpenStudy (campbell_st):

simplify the expression cos(3x)(cos(4x) - sin(4x)) then use cos(4x) =8cos^4(x) - 8cos^2(x) +1 sin(4x) = 4sin(x)cos(x) -8sin^3(x)cos(x) cos(3x) = 4cos^3(x) - 3cos(x)

OpenStudy (phi):

@open How are you getting numbers? If you were asked what is the sin(x) what would you answer? btw, are you positive you typed the question correctly?

OpenStudy (anonymous):

Yeah the question is typeed correctly

OpenStudy (campbell_st):

and cos(4x + 3x) = cos(4x)cos(3x) - sin(4x)sin(3x) and not the question asked

OpenStudy (anonymous):

Youre right, its cos at the end, not sin

OpenStudy (phi):

The simplest I see is to factor out cos(3x) cos(3x)(cos(4x)-sin(4x)